OpenStudy (anonymous):

A grocer wants to make a 10-pound mixture of peanuts and cashews that he can sell for $4.75 per pound. If peanuts cost $4.00 per pound and cashews cost $6.50 per pound, how many pounds of cashews should he use? 3 4 6 7 im thinking 4

OpenStudy (***[isuru]***):

nope.. if the pounds of cashews r "x" and the pounds of peanuts r "y" then the x + y = 10 'cause the total mixture is 10 pounds... x + y = 10 -----(1) then the price of pound of cashew multiply by "x" + the price of pound of peanut multiply by "y" is equal to the price of 10 pounds of mixture (6.5)x + 4y = price of 10 pounds of mixture but price of 1 pound of mixture = 4.75 so... (6.5)x/10 +4y/10 = 4.75 multiply the whole by 100 65x + 40y = 475 -----(2) now u want to find the value of "x" multiply (1) by 40 40x + 40y = 400 ----(3) (2) - (3) 25x = 75 x = 3 so.. the no, of pounds that he should use is 3 hope this will help ya!!